# Mapping Drivers of Greenness: Spatial Variable Selection for MODIS Vegetation Indices

### v1

@@ 4 Application to Vegetation Data
The western United States has experienced unprecedented ecological transformation in recent decades, with intensifying drought stress, record-breaking wildfire seasons, and rapid shifts in vegetation composition [williams2020large, abatzoglou2016impact]. Understanding which environmental factors drive vegetation and whether these relationships exhibit non-zero spatial effects is essential for predicting ecosystem responses and developing targeted management strategies.
@@
The response variable is the Enhanced Vegetation Index, which provides improved sensitivity over dense vegetation and enhanced atmospheric correction relative to traditional indices [huete2002overview]. We select 10 environmental predictors representing different aspects of vegetation-environment relationships: spectral reflectance bands (red, near-infrared (NIR), blue, mid-infrared (MIR)), ecosystem productivity measures (Gross Primary Productivity (GPP), Latent Heat Flux (LE)), observation geometry variables (view zenith angle, sun zenith angle, relative azimuth angle), and land surface characteristics (land cover type). We apply the transformation $\log(\text{EVI}+1)$ to improve normality and split the data into 80% training and 20% testing sets. All predictors are standardized to the $[0,1]$ range.
@@
To evaluate variable selection capabilities, we compare BSGL with lasso and adaptive lasso. Table 3 summarizes variable selection results across methods. Standard lasso selects 9 of 10 predictors, demonstrating limited selectivity by retaining nearly all variables. Adaptive lasso improves variable selection by identifying 6 variables: red reflectance, NIR reflectance, blue reflectance, MIR reflectance, GPP, and LE. We note that while GGP-GAM excels at modeling spatially varying relationships [comber2024multiscale], it is fundamentally not designed for variable selection tasks.
@@ 5 Discussion
The BSGL framework also provides a flexible platform for future extensions. Incorporating spatiotemporal components would make it possible to examine how spatial patterns evolve under climate change, distinguishing stable from shifting relationships. Extending the framework to accommodate non-Gaussian responses would widen its applicability to species distributions, biodiversity counts, and other common ecological outcomes. While the present analysis focuses on samples up to $n\sim 10,000$, recent developments in distributed Bayesian computation [guhaniyogi2022distributed, guhaniyogi2023distributed] and data sketching approaches [guhaniyogi2021sketching, guhaniyogi2025bayesiandatasketchingvarying] could be integrated with BSGL to ensure scalability to much larger datasets. Beyond vegetation monitoring, the methodology is well suited to digital soil mapping, biodiversity modeling, ecosystem service valuation, and climate impact assessment—settings characterized by high-dimensional spatial data where determining which variables merit spatially explicit treatment is both difficult and crucial.
[…]
Supplementary information Supplementary materials include detailed MCMC derivations, hyperparameter tuning results, convergence diagnostics, additional simulation results, and complete R code for implementing BSGL.
[…]
@@ Acknowledgements
Supplementary Materials
Detailed mathematical derivations, additional simulation results, and convergence diagnostics are provided in online supplementary materials.

### v2

@@ 4 Application to Vegetation Data
The western United States has experienced unprecedented ecological transformation in recent decades, with intensifying drought stress, record-breaking wildfire seasons, and rapid shifts in vegetation composition (Williams et al., 2020; Abatzoglou and Williams, 2016). Understanding which environmental factors drive vegetation and whether these relationships exhibit non-zero spatial effects is essential for predicting ecosystem responses and developing targeted management strategies.
@@
The response variable is the Enhanced Vegetation Index, which provides improved sensitivity over dense vegetation and enhanced atmospheric correction relative to traditional indices (Huete et al., 2002). We select 10 environmental predictors representing different aspects of vegetation-environment relationships: spectral reflectance bands (red, near-infrared (NIR), blue, mid-infrared (MIR)), ecosystem productivity measures (Gross Primary Productivity (GPP), Latent Heat Flux (LE)), observation geometry variables (view zenith angle, sun zenith angle, relative azimuth angle), and land surface characteristics (land cover type). We apply the transformation $\log(\text{EVI}+1)$ to improve normality and split the data into 80% training and 20% testing sets. All predictors are standardized to the $[0,1]$ range.
@@
To evaluate variable selection capabilities, we compare BSGL with lasso and adaptive lasso. Table 3 summarizes variable selection results across methods. Standard lasso selects 9 of 10 predictors, demonstrating limited selectivity by retaining nearly all variables. Adaptive lasso improves variable selection by identifying 6 variables: red reflectance, NIR reflectance, blue reflectance, MIR reflectance, GPP, and LE. We note that while GGP-GAM excels at modeling spatially varying relationships (Comber et al., 2024), it is fundamentally not designed for variable selection tasks.
@@ 5 Discussion
The BSGL framework also provides a flexible platform for future extensions. Incorporating spatiotemporal components would make it possible to examine how spatial patterns evolve under climate change, distinguishing stable from shifting relationships. Extending the framework to accommodate non-Gaussian responses would widen its applicability to species distributions, biodiversity counts, and other common ecological outcomes. While the present analysis focuses on samples up to $n\sim 10,000$, recent developments in distributed Bayesian computation (Guhaniyogi et al., 2022, 2023) and data sketching approaches (Guhaniyogi and Scheffler, 2021; Guhaniyogi et al., 2025) could be integrated with BSGL to ensure scalability to much larger datasets. Beyond vegetation monitoring, the methodology is well suited to digital soil mapping, biodiversity modeling, ecosystem service valuation, and climate impact assessment—settings characterized by high-dimensional spatial data where determining which variables merit spatially explicit treatment is both difficult and crucial.
Supplementary Information
Supplementary materials include detailed MCMC derivations, hyperparameter tuning results, convergence diagnostics, additional simulation results, and complete R code for implementing BSGL.